# Pullback of arithmetic theta series and its modularity for unitary Shimura curves

### v1

@@ 7.4. Arithmetic theta function with Bruinier Green functions I—Ehlen and Sankaran’s normalization
Now we come back to the notation and restriction of Section 7.2. The harmonic functions $F_{m}$ in Section 7.3 are now associated to the lattice $L$ with quadratic form $Q(x)=(x,x)$.
@@
Here $c_{m}^{+}(0)$ is defined in Lemma 7.7 and $\sigma_{1}(0)=-\frac{1}{24}$.
@@
(7.12) $\mathcal{Z}_{ES}^{\mathrm{tot}}(m)=\mathcal{Z}^{*}(m)+\sum_{\Phi\in\mathrm{Cusps}([[L]])}\eta_{\Phi}(m)[\mathcal{S}^{*}(\Phi)]\in{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*}),$
@@
(7.13) $\mathcal{Z}_{ES}^{\mathrm{tot}}(0)=\widehat{\Omega}^{-1}+(0,-\log D)+\sum_{\Phi\in\mathrm{Cusps}([[L]])}\eta_{\Phi}(0)[\mathcal{S}^{*}(\Phi)]\in{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*}).$
@@
(7.14) ^$\widehat{\mathbb{Z}}_{ES}^{\mathrm{tot}}(m)=(\mathcal{Z}_{ES}^{\mathrm{tot}}(m),G_{ES}(m))\in\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*}).$
Here the Ehlen-Sankaran Green function $G_{ES}(m)=G_{ES}(m,0)$ is the Green function (7.10) restricting to $\mathcal{D}\times U(V)(\mathbb{A}_{f})$, under the natural embedding $U(V)\subset\mathrm{SO}(V)$.
@@
$\Theta_{ES}(\tau)=\left(E_{2}(\tau)+\sum_{m\in\mathbb{Z}_{\geq 0}}c_{m}^{+}(0)q^{m}\right)\left(\sum_{\Phi\in\mathrm{Cusps}([[L]])}[B_{\Phi}]\right)+\Theta^{\mathrm{geo}}(\tau)$
@@
$\Theta^{\mathrm{geo}}(\tau)=-\Omega-\frac{1}{8\pi v}\sum_{\Phi\in\mathrm{Cusps}([[L]])}[B_{\Phi}]+\sum_{m\in\mathbb{Z}_{>0}}Z^{*}(m)q^{m}.$
Embedding $U(V)$ to $\hbox{SO}(V)$ and applying [EGT23, Theorem 1.2], we see that $\Theta^{\mathrm{geo}}(\tau)$ is a modular form for $\Gamma_{0}(D)$ of weight $2$ (trivial character) with values in $\operatorname{CH}_{\mathbb{C}}^{1}(S^{*})$. Now the corollary is clear. ∎
@@ 7.5. Arithmetic theta function with Bruinier Green functions II-Bruinier’s normalization
For each $m\geq 0$, there is a unique harmonic Maass (scalar valued) modular form $f_{m}$ for $\Gamma_{0}(D)$ of weight 00 such that
@@
$\tilde{f}_{m}=\sum_{\gamma\in\Gamma_{0}(D)\backslash\operatorname{SL}_{2}(\mathbb{Z})}f(\gamma\tau)\rho_{L}(\gamma^{-1})\phi_{0}$
be the associated harmonic Maass modular form for $\operatorname{SL}_{2}(\mathbb{Z})$ of weight 00 and representation $\rho_{L}$. By [${}^{+}$20, Proposition 6.1.2], we also have
@@
[…]
be the regularized theta lifting of $\tilde{f}_{m}$ as in (7.10). Then similar to [${}^{+}$20, Section 7.3], we have an arithmetic divisor
(7.18) $\widehat{\mathcal{Z}}_{B}^{\mathrm{tot}}(m)=(\mathcal{Z}_{B}^{\mathrm{tot}}(m),G_{B}(m))\in\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$
with
(7.19) $\mathcal{Z}_{B}^{\mathrm{tot}}(m)=\mathcal{Z}^{*}(m)-2\sigma_{1}(m)\sum_{\Phi\in\hbox{Cusps}[[L]]}[\mathcal{S}^{*}(\Phi)]+\begin{cases}0&\mathrm{if}m>0,\\ \widehat{\Omega}^{-1}+(0,-\log D)&\mathrm{if}m=0.\end{cases}$
@@
The issue will be stated at the end. The idea is to prove
(7.21) $\widehat{\Theta}_{B}(\tau)-\widehat{\Theta}_{ES}(\tau)=\sum_{m\geq 0}\left(c_{m}^{+}(0)\sum_{\Phi}\mathcal{S}^{*}(\Phi),-\log\|\Psi_{m}\|^{2}\right)q^{m}$
is modular. Here $\Psi_{m}$ is the Borcherds lifting of the vector valued weakly holomorphic modular form $\tilde{f}_{m}-F_{m,0}$ and has weight 00. Over $\mathbb{C}$, we know $\hbox{Div}(\Psi_{m})=c_{m}^{+}(0)\sum_{\Phi}\mathcal{S}^{*}(\Phi)$. By Corollary 7.10, we see that $\sum c_{m}^{+}(0)q^{m}$ is a modular form of weight $2$ for $\Gamma_{0}(D)$. Now the same proof as that in [ES18, Theorem 4.13] gives the modularity of (7.21).

### v2

@@ 7.4. Arithmetic theta function with Bruinier Green functions I—Ehlen and Sankaran’s normalization
Now we come back to the notation and assumptions of Section 7.2. The harmonic functions $F_{m}$ in Section 7.3 are now associated to the lattice $L$ with quadratic form $Q(x)=(x,x)$.
@@
Here $c_{m}^{+}(0)$ is defined in Lemma 7.7 and $\sigma_{1}(0)=-\frac{1}{24}$. Note that $\eta_{\Phi}$ does not depend on the choice of the cusp $\Phi$.
@@
(7.12) $\mathcal{Z}_{ES}^{\mathrm{tot}}(m)=\mathcal{Z}^{*}(m)+\sum_{\Phi\in\mathrm{Cusp}([[L]])}\eta_{\Phi}(m)[\mathcal{S}^{*}(\Phi)]\in{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*}),$
@@
(7.13) $\mathcal{Z}_{ES}^{\mathrm{tot}}(0)=\widehat{\Omega}^{-1}+(0,-\log D)+\sum_{\Phi\in\mathrm{Cusp}([[L]])}\eta_{\Phi}(0)[\mathcal{S}^{*}(\Phi)]\in{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*}).$
@@
(7.14) $\widehat{\mathcal{Z}}_{ES}^{\mathrm{tot}}(m)=(\mathcal{Z}_{ES}^{\mathrm{tot}}(m),G_{ES}(m))\in\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*}).$
Here the Ehlen-Sankaran Green function $G_{ES}(m)=G_{ES}(m,0)$ is the Green function (7.10) restricted to $\mathcal{D}\times U(V)(\mathbb{A}_{f})$, via the embedding
$\mathcal{D}\times U(V)(\mathbb{A}_{f})\rightarrow\mathcal{D}^{o}\times\mathrm{SO}(V)(\mathbb{A}_{f}),$
associated to the natural embedding $U(V)\subset\mathrm{SO}(V)$.
@@
$\Theta_{ES}(\tau)=\left(E_{2}(\tau)+\sum_{m\in\mathbb{Z}_{\geq 0}}c_{m}^{+}(0)q^{m}\right)\left(\sum_{\Phi\in\mathrm{Cusp}([[L]])}[S^{*}(\Phi)]\right)+\Theta^{\mathrm{geo}}(\tau)$
@@
$\Theta^{\mathrm{geo}}(\tau)=-\Omega-\frac{1}{8\pi v}\sum_{\Phi\in\mathrm{Cusp}([[L]])}[S^{*}(\Phi)]+\sum_{m\in\mathbb{Z}_{>0}}Z^{*}(m)q^{m}.$
Embedding $U(V)$ to $\hbox{SO}(V)$ and applying [EGT23, Theorem 1.2], we see that $\Theta^{\mathrm{geo}}(\tau)$ is a modular form for $\Gamma_{0}(D)$ of weight $2$ (trivial character) with values in $\operatorname{CH}_{\mathbb{R}}^{1}(S^{*})$. Notice $\left(\sum_{\Phi\in\mathrm{Cusp}([[L]])}[S^{*}(\Phi)]\right)$ is non-trivial as its degree is bigger than 0. Now the corollary is clear. ∎
@@ 7.5. Arithmetic theta function with Bruinier Green functions II-Bruinier’s normalization
Similar to Lemma 7.7 and [BF04, Proposition 3.11]), for each $m\geq 0$, there is a unique harmonic Maass (scalar valued) modular form $f_{m}$ for $\Gamma_{0}(D)$ of weight 0 for each $m\geq 0$ such that
@@
Indeed, the existence is essentially given by [BF04, Proposition 3.11]) without Condition (2). Two different ones differ by a constant. Condition (2) makes it unique. Lemma 7.7 (2) is another way to normalize the weight 0 harmonic Maass forms $f_{m,\mu}$.
@@
[…]
be the associated harmonic Maass modular form for $\operatorname{SL}_{2}(\mathbb{Z})$ of weight 0 and representation $\rho_{L}$. By [BHK+20, Proposition 6.1.2], we also have
@@
(7.17) $G_{B}(m)=\int^{reg}\tilde{f}_{m}\theta_{L}d\mu(\tau)$
be the regularized theta lifting of $\tilde{f}_{m}$ as in (7.10). According to [BHY15, Corollary 4.12], $G_{B}(m)$ is a Green function for
$Z^{*}(m)-2\sigma_{1}(m)\sum_{\Phi\in\hbox{Cusp}[[L]]}[S^{*}(\Phi)]$
Similar to [BHK+20, Section 7.3], we define an arithmetic divisor
(7.18) $\widehat{\mathcal{Z}}_{B}^{\mathrm{tot}}(m)=(\mathcal{Z}^{*}(m)-2\sigma_{1}(m)\sum_{\Phi\in\hbox{Cusp}[[L]]}[\mathcal{S}^{*}(\Phi)],G_{B}(m))+\begin{cases}0&\mathrm{if}\,m>0,\\ \widehat{\Omega}^{-1}+(0,-\log D)&\mathrm{if}\,m=0.\end{cases}$
@@
It suffices to prove
(7.20) $\widehat{\Theta}_{B}(\tau)-\widehat{\Theta}_{ES}(\tau)=\sum_{m\geq 0}\left(c_{m}^{+}(0)\sum_{\Phi}\mathcal{S}^{*}(\Phi),\Phi_{m}\right)q^{m}$
is modular. Here $\Phi_{m}$ is the regularized theta lifting of $\tilde{f}_{m}-F_{m,0}\in M_{0}(\rho_{L})$, and is a Green function of $c_{m}^{+}(0)\sum_{\Phi}\mathcal{S}^{*}(\Phi)$. By Borcherd’s well-known theorem theorem on Borcherds product ([Bor98]), $\Phi_{m}=-\log\|\Psi_{m}\|^{2}$, where $\Psi_{m}$ is the Borcherds lifting of $\tilde{f}_{m}-F_{m,0}\in M_{0}(\rho_{L})$ and has Borcherds product expansion. This implies that $\Psi_{m}$ is a rational section of $\omega_{\mathbb{Q}}^{cc_{m}^{+}(0)}$ for some positive rational number $c>0$. The $q$-expansion principal implies that it extends to a rational section of $\omega^{cc_{m}^{+}(0)}$, and the right hand side of (7.20) becomes
$c\sum c_{m}^{+}(0)q^{m}\hat{\omega}$
which is modular by Corollary 7.10. ∎